OpenStudy (anonymous):

1/10(a+2)=1/4(2-a) solve the equation

OpenStudy (anonymous):

a=-8/7

OpenStudy (anonymous):

a=8/7 is not correct.

OpenStudy (mertsj):

1. Multiply both sides by 20 2. Use the distributive property on both sides. 3. Add 5a to both sides 4. Subtract 4 from both sides 5. Divide both sides by 7

OpenStudy (anonymous):

\[\frac{ 1 }{ 10(a+2) } =\frac{ 1 }{ 4(2-a }\] 4(2-a)=10(a+2) 8-4a=10a+20 -16=14a a=-16/14 or -8/7

OpenStudy (mertsj):

@roza93 8-20=-12

OpenStudy (anonymous):

mertsjr, would the answer be a=2?

OpenStudy (anonymous):

-6/7

OpenStudy (mertsj):

I am not sure what the problem is. Is it this:

OpenStudy (mertsj):

\[\frac{1}{10}(a+2)=\frac{1}{4}(2-a)\]

OpenStudy (mertsj):

Waiting 30 more seconds for a response....

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i tried a different problem like it and got it right that time.